Solution (source code)

= Solution

The negative-test <probability> is $P(-)=1-0.05094=0.94906$. Applying <Bayes' theorem> again,
$$
\boxed{P(G\mid-)=\frac{P(-\mid G)P(G)}{P(-)}=\frac{0.01(0.001)}{0.94906}=1.05\times10^{-5}\quad\text{(3 s.f.)}.}
$$